OpenStudy (anonymous):

Find the derivative of f(x) = -8/x at x = 11.

OpenStudy (anonymous):

by hand or using a shortcut?

OpenStudy (kainui):

You can rewrite \[\frac{1}{x} =x^{-1}\] this is just a consequence of exponent rules lining up. Since \[x^3*x^2=(x*x*x)(x*x)=x^5\] we know exponents add. So if you have \[\frac{x^5}{x}=x^4 \] then we could just as well write this as \[x^5*x^{-1}=x^4\] by following these same rules... But the whole point of understanding that is just because the power rule works just the same here as well. \[\frac{d}{dx}(x^p)=p*x^{p-1}\]

OpenStudy (anonymous):

if you get to use a shortcut, then \[\frac{d}{dx}[\frac{1}{x}]=-\frac{1}{x^2}\]

OpenStudy (kainui):

I just assumed you weren't using the limit definition, @satellite73 has a good point.

OpenStudy (anonymous):

if you have to compute a limit by hand, then the work is different

OpenStudy (kainui):

Although not a good point in his second post where he just gives you the answer for free with no explanation.

OpenStudy (anonymous):

8/121?